# Writing Style Matters: An Examination of Bias and Fairness in Information Retrieval Systems

### v1

@@ Abstract.
The rapid advancement of Language Model technologies has opened new opportunities, but also introduced new challenges related to bias and fairness. This paper explores the uncharted territory of potential biases in state-of-the-art universal text embedding models towards specific document and query writing styles within Information Retrieval (IR) systems. Our investigation reveals that different embedding models exhibit different preferences of document writing style, while more informal and emotive styles are less favored by most embedding models. In terms of query writing styles, many embedding models tend to match the style of the query with the style of the retrieved documents, but some show a consistent preference for specific styles. Text embedding models fine-tuned on synthetic data generated by LLMs display a consistent preference for certain style of generated data. These biases in text embedding based IR systems can inadvertently silence or marginalize certain communication styles, thereby posing a significant threat to fairness in information retrieval. Finally, we also compare the answer styles of Retrieval Augmented Generation (RAG) systems based on different LLMs and find out that most text embedding models are biased towards LLM’s answer styles when used as evaluation metrics for answer correctness. This study sheds light on the critical issue of writing style based bias in IR systems, offering valuable insights for the development of more fair and robust models.
@@ 1. Introduction
Information Retrieval (IR) systems are pivotal in managing the current informational overflow, empowering users to obtain results more efficiently (Dai et al., 2024). The relevance of IR in everyday life is profound due to its use cases in diverse applications such as online web browsing, question-answering systems, digital assistants, etc. The central goal of IR systems is to identify and fetch data relevant to the user’s query (often based on their embedding similarity scores) (Hambarde and Proenca, 2023). Text embeddings for information retrieval has undergone substantial transformations throughout the past decades, from the early phase of count-based sparse embeddings such as Term Frequency-Inverse Document Frequency (TF-IDF) (Manning et al., 2008), transitioning into the second phase characterized by static dense word embeddings, including Word2Vec (Mikolov et al., 2013) and FastText (Bojanowski et al., 2017). The field then evolved into the third phase featuring contextualized embeddings like Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2018) and Embeddings From Language Models (ELMo) (Neumann et al., 2018), leading to the latest developments in the fourth phase encompassing universal text embeddings, such as General-purpose Text Embedding model (GTE) (Li et al., 2023), Beijing Academy of Artificial Intelligence General Embedding (BGE) (Xiao et al., 2023), Large Language Model (LLM) based universal text embedding (Wang et al., 2023), among others (Cao, 2024).
@@
With the widespread applications and use cases of ChatGPT, the remarkable capacities of LLMs have been highlighted, especially in terms of conversational capabilities, instruction-following and in-context learning (Cao, 2024), which has fundamentally revolutionized Information Retrieval. However, several critical issues of LLMs have also been identified including lacking of some domain knowledge and updated knowledge, and hallucinations. One solution to mitigate these limitations is Retrieval Augmented Generation (RAG) which uses text embedding to provide relevant contexts for LLMs to generate answers beyond their initial training data (Huang and Huang, 2024). LLMs are also used to generate synthetic data to fine-tune text embedding models for IR systems (Wang et al., 2023; Lee et al., 2024). The fast adoption of LLMs in IR has also brought new challenges related to bias and fairness. These issues can undermine the reliability and robustness of IR systems and may inadvertently contribute to broader societal problems (Dai et al., 2024).
@@
Recently, considerable attention has been devoted to the bias and fairness related to LLM-based applications (Dai et al., 2024). However, only a limited number of research initiatives have been concentrated on the potential bias present in text embedding models utilized for IR systems. Most existing studies ignore the fact that both human written documents and LLM-generated synthetic documents can have diverse communication or writing styles. These variations in writing styles can introduce bias and unfairness into information retrieval systems in several ways. Firstly, if the text embedding models have been trained mainly on data in a particular writing style, they may be biased and struggle to accurately interpret and retrieve information from documents written in a different style. Secondly, the system may exhibit unfairness across different writing styles. For instance, if the system has been trained to associate higher relevance with documents that use complex sentence structures, it may favor such documents in the search results even if they are not the most relevant to the user’s query while discriminating against documents in other writing styles. Finally, the writing style of queries could also introduce bias. If the system is trained with queries based on specific writing styles, it may fail to answer queries written in other styles.
@@
In this study, we fill the gap of the literature by examining the influence of varying writing styles on state of the art text embedding models with the following key contributions::
@@
  • A novel method of analyzing and quantifying fairness in text embedding models is proposed, which reveals different writing style bias in different embedding models, as well as the impact of query styles on model fairness.
@@ 2. Related Works
Bias in Information Retrieval systems: The issue of bias and unfairness in the filed of Artificial Intelligence (AI) has been a subject of extensive research in past years. The recent incorporation of LLMs into information retrieval systems has augmented the complexity of identifying bias. This is primarily because LLMs are used in multiple stages in IR systems including synthetic data generation to augment the training data of IR systems, synthetic web content generation (input of IR systems), LLMs as backbone of text embedding models, LLMs for re-ranking, LLMs as IR models and LLMs as judge to evaluate the IR system performance (Dai et al., 2024; Lee et al., 2024; Adlakha et al., 2023).
@@
In both classic and LLM-based IR systems, text embeddings play an important role (Cao, 2024), while only a few studies have focused on the potential bias and fairness in text embedding models for information retrieval. Recent research conducted by (Chen et al., 2024b) explored the immediate and long-term impacts of text generated by LLMs on the performance of RAG systems. The findings suggest that the LLM-created content consistently excels in search rankings over human-written content, which consequently reduces the visibility and influence of human-generated online contributions. In (Dai et al., 2023), the authors use LLMs to generate rewritten text copies of human written texts while preserving the same semantic meaning. They subsequently noted that neural text embedding models are biased towards LLM-generated texts.
Universal text embeddings: The rise in the popularity of text embedding techniques is evident across both industrial and academic sectors, a trend driven by their critical role in a variety of natural language processing tasks including information retrieval, question answering, semantic textual similarity, item recommendation, etc (Cao, 2024). It has been long-standing goal in the filed of text embeddings to develop a universal model which is able to address a multitude of input text length, downstream tasks, domains and languages (Li et al., 2023). With the increasing number and improved quality of diverse text datasets across different tasks, synthetic data generation, LLMs as backbones for embedding model as well as benchmarks with the focus on novel task and domain generalization such as the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022), significant advancements have been achieved in the domain of universal text embeddings recently (Cao, 2024).
@@ 5.2. Experimental results
The correctness score based on different text embedding models are shown in Figure 7 (BM25 is excluded as its values are unbounded): multilingual-e5-large-instruct shown in black color in Figure 7 (c) demonstrates the most fairness across different answer styles (with 0.03 std). However, most models show a tendency to assign higher correctness scores to FlanT5-xxl’s answers, while giving lower scores to Llama2-7B and GPT3.5-turbo. This is similar to the patterns observed in the Bleu and RougeL scores shown in Figure 6. These results indicate a prevalent bias in text embedding models towards LLM’s answer styles when used as evaluation metrics for answer correctness.
@@ Appendix A APPENDIX
Figure 8. The impact of query writing styles on the IR fairness: X-axis represents 9 different document writing styles, Y-axis shows the average ranking. Query written in style 0, 1 and 2 are shown in color red, black and orange. Dashed line is the average ranking between query and the original human written document.
Figure 9. The impact of query writing styles on the IR fairness: X-axis represents 9 different document writing styles, Y-axis shows the average ranking. Query written in style 4, 6 and 8 are shown in color red, black and orange. Dashed line is the average ranking between query and the original human written document.

### v2

@@ Abstract.
The rapid advancement of Language Model technologies has opened new opportunities, but also introduced new challenges related to bias and fairness. This paper explores the uncharted territory of potential biases in state-of-the-art universal text embedding models towards specific document and query writing styles within Information Retrieval (IR) systems. Our investigation reveals that different embedding models exhibit different preferences for document writing style, while more informal and emotive styles are less favored by most embedding models. In terms of query writing styles, many embedding models tend to match the query style with the retrieved document style, but some show a consistent preference for specific styles. Text embedding models fine-tuned on synthetic data generated by LLMs display a consistent preference for certain style of generated data. These biases in text embedding based IR systems can inadvertently silence or marginalize certain communication styles, thereby posing a significant threat to fairness in information retrieval. Finally, we also compare the answer styles of Retrieval Augmented Generation (RAG) systems based on different LLMs and find out that most text embedding models are biased towards LLM’s answer styles when used as evaluation metrics for answer correctness. This study sheds light on the critical issue of writing style based bias in IR systems, offering valuable insights for the development of more fair and robust models.
@@ 1. Introduction
Information Retrieval (IR) systems are pivotal in managing the current informational overflow, empowering users to obtain results more efficiently (Dai et al., 2024). The relevance of IR in everyday life is profound due to its use cases in diverse applications such as online web browsing, question-answering systems, digital assistants, etc. The central goal of IR systems is to identify and fetch data relevant to the user’s query (often based on their embedding similarity scores) (Hambarde and Proenca, 2023). Text embeddings for information retrieval have undergone substantial transformations throughout the past decades, from the early phase of count-based sparse embeddings such as Term Frequency-Inverse Document Frequency (TF-IDF) (Manning et al., 2008), transitioning into the second phase characterized by static dense word embeddings, including Word2Vec (Mikolov et al., 2013) and FastText (Bojanowski et al., 2017). The field then evolved into the third phase featuring contextualized embeddings like Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2018) and Embeddings From Language Models (ELMo) (Neumann et al., 2018), leading to the latest developments in the fourth phase encompassing universal text embeddings, such as General-purpose Text Embedding model (GTE) (Li et al., 2023), Beijing Academy of Artificial Intelligence General Embedding (BGE) (Xiao et al., 2023), Large Language Model (LLM) based universal text embedding (Wang et al., 2023), among others (Cao, 2024).
@@
With the widespread applications and use cases of ChatGPT, the remarkable capacities of LLMs have been highlighted, especially in terms of conversational capabilities, instruction-following and in-context learning (Cao, 2024), which has fundamentally revolutionized Information Retrieval. However, several critical issues of LLMs have also been identified, including lack of some domain knowledge and updated knowledge, and hallucinations. One solution to mitigate these limitations is Retrieval Augmented Generation (RAG) which uses text embedding to provide relevant contexts for LLMs to generate answers beyond their initial training data (Huang and Huang, 2024). LLMs are also used to generate synthetic data to fine-tune text embedding models for IR systems (Wang et al., 2023; Lee et al., 2024). The fast adoption of LLMs in IR has also brought new challenges related to bias and fairness. These issues can undermine the reliability and robustness of IR systems and may inadvertently contribute to broader societal problems (Dai et al., 2024).
@@
Recently, considerable attention has been devoted to the bias and fairness related to machine learning models (Dai et al., 2024; Cao et al., 2023). However, only a limited number of research initiatives have been concentrated on the potential bias present in text embedding models utilized for IR systems. Most existing studies ignore the fact that both human written documents and LLM-generated synthetic documents can have diverse communication or writing styles. These variations in writing styles can introduce bias and unfairness into information retrieval systems in several ways. Firstly, if the text embedding models have been trained mainly on data in a particular writing style, they may be biased and struggle to accurately interpret and retrieve information from documents written in a different style. Secondly, the system may exhibit unfairness across different writing styles. For instance, if the system has been trained to associate higher relevance with documents that use complex sentence structures, it may favor such documents in the search results even if they are not the most relevant to the user’s query while discriminating against documents in other writing styles. Finally, the writing style of queries could also introduce bias. If the system is trained with queries based on specific writing styles, it may fail to answer queries written in other styles.
@@
In this study, we fill the gap of the literature by examining the influence of varying writing styles on state of the art text embedding models with the following key contributions:
@@
  • A novel method of analyzing and quantifying fairness in text embedding models is proposed, which reveals different writing style biases in different embedding models, as well as the impact of query styles on model fairness.
@@ 2. Related Works
Bias in Information Retrieval systems: The issue of bias and unfairness in the field of Artificial Intelligence (AI) has been a subject of extensive research in past years. The recent incorporation of LLMs into information retrieval systems has augmented the complexity of identifying bias. This is primarily because LLMs are used in multiple stages in IR systems including synthetic data generation to augment the training data of IR systems, synthetic web content generation (input of IR systems), LLMs as backbone of text embedding models, LLMs for re-ranking, LLMs as IR models and LLMs as judge to evaluate the IR system performance (Dai et al., 2024; Lee et al., 2024; Adlakha et al., 2023).
@@
In both classic and LLM-based IR systems, text embeddings play an important role (Cao, 2024), while only a few studies have focused on the potential bias and fairness in text embedding models for information retrieval. Recent research conducted by (Chen et al., 2024b) explored the immediate and long-term impacts of texts generated by LLMs on the performance of RAG systems. The findings suggest that the LLM-created content consistently excels in search rankings over human-written content, which consequently reduces the visibility and influence of human-generated online contributions. In (Dai et al., 2023), the authors use LLMs to generate rewritten text copies of human written texts while preserving the same semantic meaning. They subsequently noted that neural text embedding models are biased towards LLM-generated texts.
Universal text embeddings: The rise in the popularity of text embedding techniques is evident across both industrial and academic sectors, a trend driven by their critical role in a variety of natural language processing tasks including information retrieval, question answering, semantic textual similarity, item recommendation, etc (Cao, 2024). It has been long-standing goal in the field of text embeddings to develop a universal model which is able to address a multitude of input text length, downstream tasks, domains and languages (Li et al., 2023). With the increasing number and improved quality of diverse text datasets across different tasks, synthetic data generation, LLMs as backbones for embedding model as well as benchmarks with the focus on novel tasks and domain generalization such as the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022), significant advancements have been achieved in the domain of universal text embeddings recently (Cao, 2024).
@@ 5.2. Experimental results
The correctness scores based on different text embedding models are shown in Figure 7 (BM25 is excluded as its values are unbounded): multilingual-e5-large-instruct shown in black color in Figure 7 (c) demonstrates the most fairness across different answer styles (with 0.03 std). However, most models show a tendency to assign higher correctness scores to FlanT5-xxl’s answers, while giving lower scores to Llama2-7B and GPT3.5-turbo. This is similar to the patterns observed in the Bleu and RougeL scores shown in Figure 6. These results indicate a prevalent bias in text embedding models towards LLM’s answer styles when used as evaluation metrics for answer correctness.
@@ Appendix A APPENDIX
Figure 8. The impact of query writing styles on the IR fairness: X-axis represents 9 different query writing styles, Y-axis shows the average ranking. Query written in style 0, 1 and 2 are shown in color red, black and orange. Dashed line is the average ranking between query and the original human written document.
Figure 9. The impact of query writing styles on the IR fairness: X-axis represents 9 different query writing styles, Y-axis shows the average ranking. Query written in style 4, 6 and 8 are shown in color red, black and orange. Dashed line is the average ranking between query and the original human written document.